This one involves solving by completing the square: x(squared)-6x+6=0
I came up with an answer of 3 is this correct? Here is my work simplified
x(squared)-6x=-6
x(squared)-6x+9=9-6
x(squared)+3=+3
then (x-3)squared = 3
x=3

Answer by tutorcecilia(2152)

Just complete by taking the square root of both sides:
(x-3)^2=3
sqrt(x-3)^2=sqrt3
x-3=sqrt3
x-3+3 = sqrt3 + 3
x=3+sqrt3